Find the x & y to make this equation true
3x-4+(y-3)i=11+5i

Answer by lwsshak3(11628)

Find the x & y to make this equation true
3x-4+(y-3)i=11+5i
3x-4=11
3x=15
x=5
..
y-3=5
y=8
check:
(3*5-4)+(8-3)i=11+5i